OpenStudy (anonymous):

Please Show All Work: Solve for g. (g – 65) × 19 = 266 g = Please Show All Work: Solve for g. (g – 65) × 19 = 266 g = @Mathematics

OpenStudy (anonymous):

19g - 1235 = 266 19g = 1501 g = 79

OpenStudy (anonymous):

19g-65X19=266(distributive property)

OpenStudy (anonymous):

19g-1235=266

OpenStudy (anonymous):

adding 1235 on both sides we get

OpenStudy (anonymous):

19g=266+1235

OpenStudy (anonymous):

19g=1501

OpenStudy (anonymous):

dividing both sides by 19 we get g=79